OpenStudy (anonymous):

Convert (5, 3π/4) into rectangular coordinates.

jimthompson5910 (jim_thompson5910):

You are given r = 5 theta = 3pi/4

jimthompson5910 (jim_thompson5910):

you want a point in the form (x,y) or in the form x+y*i use the formulas x = r*cos(theta) y = r*sin(theta) to get the values of x and y

jimthompson5910 (jim_thompson5910):

theta is already given, theta = 3pi/4 this is assuming (5, 3π/4) is in polar form (r,theta)

OpenStudy (jhannybean):

Cylindrical \(\rightarrow\) Rectangular \(r ~ , \theta ~ , z\) \(\rightarrow\) \(x ~ , ~ y ~ , ~ z\)

OpenStudy (jhannybean):

\[\sf x = r\cos(\theta) ~, y = r\sin(\theta) ~ , z =z\]

OpenStudy (dan815):